OpenStudy (anonymous):

4. As a valet at the mall, Davis works 5-hour shifts. He earns $7 per hour and keeps 90% of his tip money. How much does Davis need to be tipped so that he takes home exactly $62.50, before taxes are deducted, per shift? $275.00 $27.50 $30.56 $69.44 i need help

OpenStudy (gorv):

in 5 hour earning=7*5=35 let tip=x tip earn=x*90/100 total=35+x*90/100=62.5 subtracting 35from both side 90x/100=62.5-35=27.5 multiplying both side by 100 90x=2750 divide both side by 90 x=2750/90=275/9

OpenStudy (gorv):

=30.56

OpenStudy (anonymous):

Thank You